OpenStudy (anonymous):

Explain the conditions necessary to multiply two matrices, each of a different order. Provide an example with your explanation.

OpenStudy (anonymous):

A_mxn * B_nxp: pay attention at the size of matrices A and B. the number of column of A must be the same with the number of row of B

OpenStudy (anonymous):

you can make up any example by yourself, just follow that rule.

OpenStudy (anonymous):

\[\left[\begin{matrix}3 & 4 &5\\ 6 & 7 & 8\end{matrix}\right]* \left[\begin{matrix}9 &6 \\ 4 & 5\\7 & 3\end{matrix}\right]\]